Solve the given initial-value problem. d2y dθ2 + y = 0, y(π/3) = 0, y'(π/3) = 6
Answer #1

y et o = 2 6²+)y=0 m+1=0 ay mati » auxilary wots - <uluър • Уг с дъ+ < sinx 9 () -0 =) Os C, CS CT )+csin() =) 0:6,()+() - 16
Answer #2

d2y dθ2 + y = 0, y (π / 3) = 0, y '(π / 3) =2
Answer #3

d2y dθ2 + y = 0, y (π / 3) = 0, y '(π / 3) =2